# DCT-CryptoNets: Scaling Private Inference in the Frequency Domain

### v1

@@ 2.4 Hyper-quantization Background
Hyper-quantization is a lossy compression technique that improves the efficiency of neural network models by reducing the precision of numerical representations (weights and activations). This not only lowers memory requirements and accelerates computation but also naturally aligns with the integer-based polynomial representation in TFHE, facilitating efficient computation within the encrypted domain. However, hyper-quantization often leads to a trade-off in accuracy as the limited representation can introduce errors.
Quantization-aware training (QAT) can help mitigate these accuracy trade-offs by simulating the effects of quantization during the training process. By constraining weights and activations to fixed-point representations, QAT intentionally introduces quantization noise during training, enabling the network to learn and adapt to these errors through gradient-based optimization. Crucially, "fake quantization" is employed to maintain full-precision weights for smoother convergence [27]. In the context of uniform quantization, we convert a real value $r\in[\alpha,\beta]$ into a b-bit integer $q=\lfloor\frac{r}{S}+Z\rfloor$, where $S=\frac{\beta-\alpha}{2^{b}-1}$ represents the quantization scale and Z is the zero-point. Dequantization is achieved by the inverse operation: $r=S(q-Z)$. Previous efforts to quantize FHENNs have produced promising results. Notably, [32] achieved successful inference on a VGG-9 network using an 8-bit quantized TFHE scheme with programmable bootstrapping (PBS). Their approach achieved 87.5% accuracy on a VGG-9 network but still had significant latency of 18,000 seconds for inference on the CIFAR-10 dataset. More aggressive quantization can also yield latency reductions, such as a 40% decrease in runtime when moving from an 8-bit to a 2-bit BFV-based FHENN for CIFAR-10 [20]. Quantization-aware training facilitates the seamless conversion of our models into a TFHE-compatible framework, enabling us to maintain high accuracy while significantly reducing computational overhead.
@@ 2.5 Frequency Domain Background
Figure 2: Frequency base pre-processing proposed by [35] and DCT-CryptoNets proposed network architecture. Changes from method to method are emphasized in darker purple. Going from RGB-based ResNet head to DCT ResNet head includes changes to first convolution layer and the exclusion of the pooling layer. Changes from the Original DCT-based ResNet head to DCT-CryptoNets ResNet head include pruning the ReLU operator after the first convolution layer as proposed by [11, 14].
In a departure from prior work, our method operates directly in the frequency domain, utilized the Discrete Cosine Transform (DCT) - a key component of JPEG compression, widely used for image transmission and storage. This allows our method to be directly compatible with machine learning as a service (MLaaS) systems and other applications where image data is transmitted in compressed formats for remote analysis. The 1-D DCT transform uses the following equation where $x_{n}$ represents a signal and $X_{k}$ represent the DCT coefficients. DCT in 2-D is the same 1-D equation applied to the width and height dimensions.
@@
While the DCT itself is lossless and reversible, by selectively retaining only low-frequency components we can reduce data requirements while preserving image classification accuracy [35]. This approach leverages the human visual system’s differential sensitivity to various frequency components, allowing models to focus on the most perceptible elements. Adversarially robust object recognition models designed to resist perceptible perturbations have been shown to primarily rely on this low-frequency information [33, 23]. Additionally, utilizing DCT components directly can improve the privacy preservation of recognition systems against various white-box and black-box attacks [34]. While previous work [25] has explored DCT for homomorphic image compression and decompression, our method uniquely retains the encrypted data in the frequency domain throughout the CNN inference process. The shift from high-spatial RGB to low-frequency DCT components in our training approach offers a dual advantage. By prioritizing perceptually relevant information, it holds the potential for improved image classification. Simultaneously, the spatial dimensionality reduction of the DCT significantly lowers the computational demands of homomorphic inference, particularly by reducing the burden of non-linear activations and homomorphic bootstrap operations.
@@ 3 Methodology
In this paper we propose DCT-CryptoNets, a novel approach for homomorphic inference that achieves scalability and low latency through a training methodology and network topology specifically designed for the frequency domain. Our approach is composed of two key components: (1) A network architecture optimized for the Discrete Cosine Transform (DCT) domain with strategic ReLU reductions and (2) Encryption of the trained network into a TFHE-compatible scheme.
3.1 Network Topology in the DCT Domain with ReLU Reductions
Table 1: Comparison of multiply and accumulate, non-linear activation, programmable bootstrap and overall homomorphic operations for RGB-based (top row) and DCT-based (middle rows) ResNet-18 network. We see that when moving from the RGB to the DCT domain there is only a slight reduction in MACs but a larger reduction in ReLUs and HOPs. We also see the same reduction in ReLU operations regardless on the number of frequency channels sub-selected, allowing us to use a DCT representation that provides the best task specific accuracy. We observe an additional 11% reduction in HOPs when applying DCT-CryptoNets to a 224$\times$224 vs. 448$\times$448 image respectively, showing that the computational effectiveness of DCT optimizations improve as image resolution increases.
@@ 3.1 Network Topology in the DCT Domain with ReLU Reductions
Following the methodology in [35], we first transform RGB images to YCrCb color space, then apply the Discrete Cosine Transform (DCT) to generate frequency domain representations. We adopt standard JPEG compression’s 8$\times$8 filtering scheme, yielding 64 frequency channels per YCrCb color space. This process effectively downsamples spatial dimensions (H, W) by a factor of 8 while increasing channel dimensions (C) by a factor of 82. Subsequently, we sub-select a fixed portion of the low-frequency channels, emphasizing luma (Y) components due to their higher perceptual relevance compared to chroma (Cr, Cb). The selection adheres to a standard 4:2:2 luma-chroma subsampling ratio. Notably, the total number of homomorphic operations (HOPs) remains relatively consistent regardless of the number of sub-selected frequency channels, allowing for flexibility in choosing a network that provides high accuracy (see Table 1). The DCT’s 8$\times$8 filtering of spatial information into frequency channels proves particularly advantageous as image size increases. We observe that DCT becomes increasingly valuable for HOP optimization as we scale to ImageNet-sized images and beyond. For 4482 images, we observe a 44.1% reduction in HOPs compared to a 33.1% reduction for 2242 images. While the percentage of ReLU operations removed remains constant, this amplified HOP reduction at larger image sizes can be attributed to the decreased need for bootstrapping operations, a direct consequence of fewer ReLU activations (see Table 1). This highlights the scalability of DCT-CryptoNets in handling larger inputs.
Given the distinct dimensionality of frequency-domain features (reduced H and W, increased C), we modify the initial stride-2 convolution layer, commonly found in CNNs, with a stride-1 layer. The subsequent layer’s channel size is then adjusted to accommodate the transformed input (see Figure 2). This approach minimally modifies existing CNN architectures to accept frequency-domain input. Further optimization involves pruning the ReLU activations following the channel-widening convolution layer. This reduction in ReLU operations does not significantly impact accuracy, as earlier-layer ReLUs typically exhibit lower importance than those in later layers [14]. While the skip connection pruning proposed in [11] also reduces ReLU count, it incurs a much greater accuracy cost and also increase HOPs (see Section A.2). Therefore, we opted for a network architecture that specifically incorporates first-layer ReLU pruning as shown in Fig. 2.
While DCT offers significant benefits for larger images, its direct application to lower resolutions can lead to excessive spatial dimension reduction. To address this challenge, we introduce a second specialized architecture, tailored for smaller inputs (e.g., 32$\times$32 pixels). This adaptation utilizes a 4$\times$4 DCT filter, resulting in 48 frequency channels, and incorporates architectural adjustments to maintain spatial resolution and prevent information loss (see Section A.3 for an architecture deep-dive). Despite an increased parameter count, this architecture achieves comparable latency and a 4% accuracy improvement over a traditional ResNet-20 on CIFAR-10. Furthermore, even with further spatial reduction from 16$\times$16 to 8$\times$8, we observe similar accuracy to a traditional ResNet-20 coupled with a 4.2$\times$ latency improvement (see Table 2). This demonstrates the effectiveness of DCT-CryptoNets even in low resolution settings.
@@
Table 2: Evaluating performance on low resolution images by comparing the traditional ResNet-20 and our proposed DCT-CryptoNets ResNet-20 on CIFAR-10. For DCT-CryptoNets ResNet-20, various combinations of retained frequency channels and spatial input sizes were explored. Our findings indicate that retaining the top 24 low-frequency channels (out of 48) and a spatial size of 16$\times$16 leads to a 4% accuracy improvement over the traditional ResNet-20 model while maintaining comparable latency.
@@
3.2 Encrypting into a TFHE Compatible Scheme
We first employ a quantization-aware training (QAT) framework on the preprocessed data and chosen network topology. To achieve an optimal balance between accuracy and latency for our FHENNs, we systematically explored various quantization levels (see Appendix A.1). Our approach employs a 4-bit symmetric quantization scheme with a fixed zero-point for all tasks except ImageNet, where 5-bit quantization is used. These choices strike a balance between efficiency and accuracy, allowing us to perform the entire neural network computation directly on encrypted numbers within the TFHE framework.
To fully encrypt the trained neural network within the TFHE scheme, we employ an automated optimization process [2]. This process balances three key hyperparameters: (1) the circuit bit-width, defined as the minimum number of bits required to represent the largest integer arising during computation. This generally mimics the 4-bit or 5-bit integer used during QAT; (2) precision rounding for model accumulators, which implements a scaling factor used to remove least significant bits; and (3) the desired error probability for programmable bootstrapping (PBS) operations. Unlike polynomial approximation (PAAs) that introduce inherent approximation errors, PBS error probability is a tunable hyperparameter. It controls the likelihood of a small error value added to the look-up table (LUT) operation. By carefully tuning these hyperparameters, the optimization process determines the cryptosystem parameters that ensure efficient execution.
@@ 3.2 Encrypting into a TFHE Compatible Scheme
Determining optimal cryptographic hyperparameters involves empirical exploration due to the indirect control over the underlying cryptosystem parameters. Decreasing the PBS error probability or increasing the precision rounding threshold to allow more bits necessitates additional homomorphic operations (HOPs). This presents a trade-off: while enhancing accuracy, it also increases latency. We notice that for simpler tasks, a precision rounding of 6 and a PBS error probability of 0.01 achieve a good balance. However, more complex tasks like ImageNet require further hyperparameter tuning to maintain accuracy. For instance, increasing precision rounding to 7 reduces bit truncation but leads to a higher number of HOPs, thus increasing latency (see Appendix A.4 for a detailed analysis).
@@ 4.1 Experimental Setup
DCT-CryptoNets utilizes several tools and frameworks. Image conversion to the DCT domain is performed using libjpeg-turbo, Brevitas [28] facilitates quantization-aware training, and Concrete-ML [36] enables encryption within the TFHE scheme. Neural network training is conducted on an RTX A40 GPU and FHENN latency measurements are obtained from dual AMD Ryzen Threadripper PRO 5965WX processors (96-threads total). Each network is trained on CIFAR-10, mini-ImageNet, Imagenette, and ImageNet, with varying configurations of DCT components. Data augmentation strategies include random cropping, color jitter, and random flipping. To assess FHENN accuracy, we leverage Concrete-ML’s simulation accuracy feature, allowing us to efficiently estimate expected accuracy and gather valuable metrics on the compiled FHENN. Latency measurements are conducted by processing a single image at a time on a dedicated machine with no background tasks running, ensuring maximum utilization of all available threads.
@@ 4.2 Accuracy and Reliability Enhancements with Low-Frequency Information
DCT-based networks consistently matched or outperformed the accuracy of their RGB counterparts, particularly on larger datasets (mini-ImageNet, ImageNet), due to their focus on perceptually relevant low-frequency information. Generally, retaining the top 64 low-frequency channels yielded the best accuracy. This suggests that lossless representations (192 channels) contain unnecessary high-frequency components that hinder learning. Thus, some degree of "lossy-ness" allows the network to prioritize visually salient information, further improving accuracy. Furthermore, DCT-based methods exhibited narrower confidence intervals, indicating greater accuracy stability in the encrypted domain. This improved reliability stems from the -30% reduction in programmable bootstrap (PBS) operations (see Table 1), which individually contribute to error accumulation. DCT-CryptoNets meets or exceeds the accuracy of RGB-based methods while also enhancing the reliability of encrypted accuracy.
4.3 Evaluation of Accuracy and Latency
Table 4: Performance comparison of DCT-CryptoNets and prior work. We selected the 64$\times$562 DCT representation for ImageNet and 24$\times$162 for CIFAR-10 due to their superior accuracy. The RGB variants, trained on 3$\times$322 or 3$\times$2242 inputs, serve as baselines without DCT optimization so demonstrate the isolated impact of DCT-based optimizations. DCT-CryptoNets show comparable performance on smaller models/datasets but excel in scaling to larger ones, highlighting its potential for real-world applications.

### v2

@@ 2.4 Hyper-quantization Background
Hyper-quantization is a technique that improves the efficiency of neural networks by reducing the precision of numerical representations (weights and activations). This not only lowers memory requirements and accelerates computation but also naturally aligns with the integer-based polynomial representation in TFHE, facilitating efficient computation within this encrypted domain. However, hyper-quantization often leads to a trade-off in accuracy as the limited representation can introduce errors.
Quantization-aware training (QAT) can help mitigate these accuracy trade-offs by simulating the effects of quantization during the training process. By constraining weights and activations to fixed-point representations, QAT intentionally introduces quantization noise during training, enabling the network to learn and adapt to these errors through gradient-based optimization (Nagel et al., 2021). In the context of uniform quantization, we convert a real value $r\in[\alpha,\beta]$ into a b-bit integer $q=\lfloor\frac{r}{S}+Z\rfloor$, where $S=\frac{\beta-\alpha}{2^{b}-1}$ represents the quantization scale and Z is the zero-point. Previous efforts to quantize FHENNs have produced promising results. Notably, Stoian et al. (2023) achieved inference on a VGG-9 network using an 8-bit quantized TFHE scheme with programmable bootstrapping (PBS). Their approach achieved 87.5% accuracy on a VGG-9 network but still had significant latency of 18,000 seconds for inference on the CIFAR-10 dataset. More aggressive quantization can also yield latency reductions, such as a 40% decrease in runtime when moving from an 8-bit to a 2-bit BFV-based FHENN for CIFAR-10 (Legiest et al., 2023). Quantization-aware training facilitates the seamless conversion of our models into a TFHE-compatible framework, enabling us to maintain high accuracy while significantly reducing computational overhead.
@@ 2.5 Frequency Domain Background
Figure 2: DCT-CryptoNets’ frequency encoding (based on Xu et al. (2020)) and ResNet-18 network architecture. Modifications from an RGB-based network to DCT-CryptoNets are emphasized in bold and darker purple. These include kernel size and downsampling of the first convolution layer, as well as exclusion of both the ReLU operator and pooling layer after the first convolutional layer. This approach requires minimal modification of existing networks to utilize DCT, making conversion for many potential applications simple.
In a departure from prior private inference work and frequency-domain techniques (Bian et al., 2020; Lou et al., 2020; Kim & Guyot, 2023), our method learns features from the frequency-domain representation of images, enabling the network to learn from the rate of change of intensities. This is done through the utilization of the Discrete Cosine Transform (DCT) - a key component of JPEG compression, widely used for image transmission and storage (Gueguen et al., 2018; Ehrlich & Davis, 2019). This allows our method to be directly compatible with machine learning as a service (MLaaS) systems and other applications where image data is transmitted in compressed formats for remote analysis (see Section A.2). The 1-D DCT transform uses the following equation where $x_{n}$ represents a signal and $X_{k}$ represent the DCT coefficients transformed via sinusoidal bases. DCT in 2-D is the same 1-D equation applied to the width and height dimensions.
@@
While DCT itself is lossless and reversible, by selectively retaining only low-frequency components we can reduce data requirements while preserving image classification accuracy (Xu et al., 2020). This approach leverages the human visual system’s differential sensitivity to various frequency components, allowing models to focus on the most perceptible elements. Adversarially robust object recognition models designed to resist perceptible perturbations have been shown to primarily rely on this low-frequency information (Subramanian et al., 2023; Luo et al., 2022). Additionally, utilizing DCT components directly can improve the privacy preservation of recognition systems against various white-box and black-box attacks (Wang et al., 2022). While previous work, such as Mertens et al. (2024), has explored DCT for homomorphic image compression and decompression, our method uniquely retains the encrypted data in the frequency domain throughout the CNN inference process. The shift from high-spatial RGB to low-frequency DCT components in our training approach offers a dual advantage. By prioritizing perceptually relevant information, it holds the potential for improved image classification. Simultaneously, the spatial dimensionality reduction of DCT significantly lowers the computational demands of homomorphic inference, particularly by reducing the burden of non-linear activations and homomorphic bootstrap operations due to the fact that non-linear activations are disproportionately concentrated in network layers with high spatial dimensionality.
@@ 3 Methodology
DCT-CryptoNets is composed of two key pieces: (1) We propose network architectures for low-frequency DCT components coupled with strategic reductions in ReLU activations for more optimal computation in the encrypted domain. These architectural designs deliver efficiency gains compared to traditional RGB-based networks across various image resolutions, even in the challenging case of low-resolution inputs where direct DCT application may be less effective. (2) We present a framework for quantization-aware training on these low-frequency DCT components, enabling seamless encryption into a TFHE-compatible scheme. We present multiple ablations into the effect of quantization and cryptographic hyperparameters on accuracy and latency in the encrypted domain.
3.1 Network Architecture in the DCT Domain with ReLU Reductions
Table 1: Comparison of multiply-accumulate (MAC), ReLU, programmable bootstrap (PBS) and total homomorphic operations (HOPs) for RGB-based and DCT-based ResNet-18 networks. DCT consistently reduces ReLUs, PBS and HOPs, even with varying numbers of retained low-frequency channels, enabling accuracy-driven selection. Notably, DCT’s HOP reduction improves with higher image resolutions (e.g., 11% greater HOP reduction for 448$\times$448 vs. 224$\times$224 images) due to additional 13.9% reduction in PBS operations, showcasing its scalability.
@@ 3.1 Network Architecture in the DCT Domain with ReLU Reductions
For encoding DCT frequency tensors from images, we first transform RGB images to the YCrCb color space, then apply 2D-DCT operations to generate frequency domain representations (see Figure 2). We adopt standard JPEG compression 8$\times$8 filtering scheme, yielding 64 frequency components per YCrCb color space. This process effectively downsamples spatial dimensions (H, W) by a factor of 8 while increasing channel dimensions (C) by a factor of 82. Subsequently, we sub-select a fixed portion of the frequency components, emphasizing luma (Y) components due to their higher perceptual relevance compared to chroma (Cr, Cb). This sub-selection of frequency components can be viewed as varying the level of "lossy-ness" of an image from a highly lossy 6-component to a loss-less 192-component representation. We demonstrate the impact of varying the level of sub-selected frequency components (6, 24, 48, 64, and 192) on the number of homomorphic operations (HOPs) and model accuracy. Notably, the total number of HOPs remains consistent regardless of the sub-selection of frequency components, allowing for flexibility in choosing an input representation that provides high accuracy (see Table 1). The DCT’s filtering of spatial information into frequency components proves particularly advantageous as image size increases. We observe that DCT becomes increasingly valuable for HOP optimization as we scale to ImageNet-sized images and beyond. For 448$\times$448 images, we observe a 44.1% reduction in HOPs compared to a 33.1% reduction for 224$\times$224 images (see Table 1). While the percentage of ReLU operations removed remains constant due to the inherent structure of DCT components, this amplified HOP reduction at larger image sizes can be attributed to the decreased need for bootstrapping operations, a direct consequence of fewer ReLU activations. This highlights the scalability of DCT-CryptoNets in handling larger inputs.
@@
Table 2: Evaluating performance on low resolution images with a traditional ResNet-20 and proposed DCT-CryptoNets ResNet-20 on CIFAR-10. DCT-CryptoNets ResNet-20, with 24 retained low-frequency channels and 16$\times$16 spatial input, achieves a 4% accuracy improvement over the baseline while maintaining latency.
@@
Given the distinct dimensionality of frequency-domain features (reduced H and W, increased C), we modify the initial stride-2 convolution layer, commonly found in CNNs, with a stride-1 layer. The subsequent layer’s channel size is then adjusted to accommodate the transformed input (see Figure 2). Further optimization involves pruning the ReLU activations following the channel-widening convolution layer. This reduction in ReLU operations does not significantly impact accuracy, as earlier-layer ReLUs typically exhibit lower importance than those in later layers (Jha et al., 2021). While the skip connection pruning introduced in Ghodsi et al. (2020) also reduces ReLU count, it incurs a much greater accuracy cost and also increase HOPs in TFHE-based networks (see Section A.3). Therefore, we opted for a network architecture that specifically incorporates first-layer ReLU pruning. Figure 2 shows both the process for encoding DCT frequency tensors as well as the network architecture needed for handling DCT tensors.
While DCT offers significant benefits for larger images, its direct application to lower resolutions can lead to excessive spatial dimension reduction. To address this challenge, we introduce a second specialized architecture, tailored for smaller inputs (e.g., 32$\times$32 pixels). This adaptation utilizes a 4$\times$4 DCT filter, resulting in 48 total frequency channels, and incorporates architectural adjustments to maintain spatial resolution and prevent information loss (see Section A.4 for an architecture deep-dive). Despite an increased parameter count due to expansion of the channel dimensions, DCT-CryptoNets achieves similar latency coupled with a 4% accuracy improvement over a traditional ResNet-20 on CIFAR-10 (DCT-CryptoNets 24$\times$162 vs. Traditional 3$\times$322 in Table 2). Furthermore, even with further spatial reduction from 16$\times$16 to 8$\times$8, we observe similar accuracy to a traditional ResNet-20 coupled with a 4.2$\times$ latency improvement. This demonstrates the effectiveness of DCT-CryptoNets, not only for high resolution images but for low resolution images as well.
3.2 Training and Encrypting into a TFHE-Compatible Scheme
DCT-CryptoNets employs a quantization-aware training (QAT) framework to learn low-frequency components. To achieve an optimal balance between accuracy and latency for our FHENNs, we systematically explored various quantization levels (see Section A.5). Our approach employs a 4-bit symmetric quantization scheme with a fixed zero-point for all tasks except ImageNet, where 5-bit quantization is used. These choices strike a balance between efficiency and accuracy, allowing us to perform the entire neural network computation directly on encrypted data within the TFHE framework.
To fully encrypt the trained neural network within the TFHE scheme, we employ an optimization process (Bergerat et al., 2023). This process balances three key cryptographic hyperparameters: (1) the circuit bit-width, defined as the minimum number of bits required to represent the largest integer arising during computation. This generally mimics the 4-bit or 5-bit integer used during QAT; (2) precision rounding for model accumulators, which implements a scaling factor used to remove least significant bits; and (3) the desired error probability for programmable bootstrapping (PBS) operations. Unlike polynomial approximation (PAAs) that introduce inherent approximation errors, PBS error probability is a tunable hyperparameter. It controls the likelihood of a small error value added to the look-up table (LUT) operation. By carefully tuning these hyperparameters, the optimization process determines the cryptosystem parameters that ensure efficient execution.
@@ 3.2 Training and Encrypting into a TFHE-Compatible Scheme
Determining optimal cryptographic hyperparameters involves empirical exploration due to the indirect control over the underlying cryptosystem parameters. Decreasing the PBS error probability or increasing the precision rounding threshold to require more bits necessitates additional homomorphic operations (HOPs). This presents a trade-off: while enhancing accuracy, it also increases latency. We notice that for simpler tasks, a precision rounding of 6 and a PBS error probability of 0.01 achieve a good balance. However, more complex tasks like ImageNet require further hyperparameter tuning to maintain accuracy. For instance, increasing precision rounding to 7 reduces bit truncation but leads to a higher number of HOPs, thus increasing latency (see Appendix A.6 for a detailed analysis).
@@ 4.1 Experimental Setup
DCT-CryptoNets utilizes several tools and frameworks. Image conversion to the frequency domain is performed using libjpeg-turbo, Brevitas (Pappalardo, 2023) facilitates quantization-aware training, and Concrete-ML (Zama, 2022) enables encryption within the TFHE scheme. Neural network training is conducted on an RTX A40 GPU and FHENN latency measurements are obtained from dual AMD Ryzen Threadripper PRO 5965WX processors (96-threads total). Each network is trained on CIFAR-10, mini-ImageNet, Imagenette, and ImageNet, with varying configurations of DCT components. To assess FHENN accuracy, we leverage Concrete-ML’s simulation accuracy feature, allowing us to efficiently estimate expected accuracy and gather valuable metrics on the compiled FHENN. Latency measurements are conducted by processing a single image at a time on a dedicated machine with no background tasks running, ensuring maximum utilization of all available threads. All training, quantization and cryptographic hyperparameters are also disclosed in Section A.7.
@@ 4.2 Accuracy and Reliability Enhancements with Low-Frequency Information
DCT-based networks consistently matched or outperformed the accuracy of their RGB counterparts, particularly on larger datasets (mini-ImageNet & ImageNet), due to their focus on perceptually relevant low-frequency information. Generally, retaining the top 64 low-frequency channels yielded the best accuracy. Suggesting that loss-less representations (192 channels) contain unnecessary high-frequency components that hinder learning. Thus, some degree of "lossy-ness" allows the network to prioritize visually salient information, further improving accuracy. Furthermore, DCT-based methods exhibited narrower confidence intervals, indicating improved inference precision in the encrypted domain. This improved reliability stems from the -30% reduction in programmable bootstrapping (PBS) (see Table 1), which contribute to error accumulation. Overall, DCT-CryptoNets meets or exceeds the accuracy of RGB-based methods while also enhancing the reliability of private inference.
4.3 Comparative Benchmarking of Accuracy and Latency
DCT-CryptoNets demonstrates competitive accuracy compared to existing methods on both CIFAR-10 and ImageNet (see Table 4). While CKKS-based methods demonstrate slightly lower latency on smaller models due to their utilization of single instruction multiple data (SIMD) packing, the error accumulation issue of polynomial approximated activation functions (PAAs) hinders CKKS-based methods from scaling to larger networks and high-resolution images like those in ImageNet. It’s crucial to note that CKKS’s focus on amortized runtime through SIMD packing, a feature unavailable in TFHE, makes direct latency comparisons via normalization less meaningful (see Section A.8 for more information on the exact compute resources used for each TFHE-based methods).
Comparing DCT-CryptoNets to SHE (Lou & Jiang, 2019), the only other published method demonstrating scalability to ImageNet, we observe comparable accuracy and latency for smaller models and images. It’s worth noting that SHE’s fastest model on CIFAR-10 is a custom CNN without skip connections, contributing to its slight latency advantage over DCT-CryptoNets’ ResNet-20 model. As network or image size increases, DCT-CryptoNets showcases superior scalability, with progressively greater latency improvements compared to SHE culminating in a 5.3$\times$ speedup on ImageNet. Furthermore, even without DCT optimizations, larger DCT-CryptoNets ResNet-18 models outperform SHE on both CIFAR-10 and ImageNet. This improvement can be attributed to SHE’s reliance on a leveled TFHE scheme, which necessitates higher multiplicative depths and larger ciphertexts leading to latency amplifications in deeper networks. In contrast, by leveraging programmable bootstrapping, DCT-CryptoNets mitigates these limitations and enables efficient handling of larger models. Isolating the impact of DCT, we compare the latency of DCT-CryptoNets trained on RGB components to DCT-CryptoNets trained on DCT components (see Table 4). Across various model and image sizes, we observe consistent latency improvements ranging from 1.7$\times$ to 2.4$\times$ when utilizing DCT, independent of training, quantization and cryptographic hyperparameters. Overall, DCT-CryptoNets achieves performance on par with prior work on smaller images and networks, while showcasing superior scalability and efficiency gains when applied to larger, more complex tasks.
Table 4: Performance benchmarking of DCT-CryptoNets. We evaluate performance using the 64$\times$562 DCT representation for ImageNet and 24$\times$162 for CIFAR-10, chosen for their superior accuracy. RGB variants of DCT-CryptoNets, trained on 3$\times$322 or 3$\times$2242 inputs, serve as baselines to isolate the impact of DCT optimizations. DCT-CryptoNets offers comparable performance on smaller models and datasets but excel in scaling to larger ones.